Seafoam:

When x=3 and y=-6, what is n? x^2-16+7*y=n

QuestionCoveBot:

So.... \[3^2-16+7*-6=n\]

QuestionCoveBot:

Do... \[3^2\]...to get 9. Then you have: \[9-16+7*-6=n\]

QuestionCoveBot:

What's 9-16+7 then?

Seafoam:

0?

QuestionCoveBot:

Yes. And when you multiply/divide by 0, the end answer is always 0. (: so then, obviously, 0*-6 is 0. Which then makes n=0 so therefore, n=0

Seafoam:

Thanks :)

QuestionCoveBot:

yur welcome- xD

QuestionCoveBot:

also congrats on purple (:

jhonyy9:

@questioncovebot wrote:
So.... \[3^2-16+7*-6=n\]
maybe more accepted (-6) in parentheses

Naughtyyy:

Don't listen to QuestionCoveBot he's actually a bot at math.

Naughtyyy:

n = 3^2 -16 + 7 * -6

QuestionCoveBot:

I already did this, and the question is answered. No need to answer it again. >.>

Naughtyyy:

7 * -6 = -42 so n = 3^2 - 16 -42

Naughtyyy:

You're wrong. Lol.

Naughtyyy:

3^2 -58

Naughtyyy:

9 - 58

Naughtyyy:

n = -49

Naughtyyy:

@questioncovebot explain to me how it's n = 0 and not n = -49

Naughtyyy:

;)

jhonyy9:

@questioncovebot wrote:
Do... \[3^2\]...to get 9. Then you have: \[9-16+7*-6=n\]
here first you need multiplie 7 by (-6) dont add (-16) to 7 you dont know PEMDAS

Naughtyyy:

1 attachment
Naughtyyy:

Actually a clown.